OpenStudy (scott208):

What is 0.8% of 125% of 8128? Express your answer as a decimal.

OpenStudy (mathmale):

In your shoes I'd find this first: "125% of 8128." How have you done similar problems in the past? Show your work.

OpenStudy (scott208):

well first i break down that 1% of 8128 is 81.28. i multiply that by 25 and get 2032. I add that to the 8128 and get 10160

OpenStudy (scott208):

WAIT! is it 81.28 is that the 0.8% of 10160?

OpenStudy (scott208):

@mathmale

OpenStudy (mathmale):

find this first: "125% of 8128" Is that percentage equal to 8128, or smaller than 8128, or larger than 8128? Try again to find "find this first: "125% of 8128."

OpenStudy (mathmale):

In your shoes I'd convert "125%" into a quantity with a decimal fraction.

OpenStudy (scott208):

125% is the 10160 and it said that 81,28 was correct

OpenStudy (mathmale):

One thing at a time, please. 125% converts to the mixed numer 1.25, and 1.25 times 8128 does come out to 10160. Good. Now find 0.8% of 10160. Please show your work...every step of it.

OpenStudy (scott208):

100%=10160 10%=1016 1%=101.6 101.6/10=10.16 (= 0.1%) 101.6-10.16=91.44(=0.9%) 91.44-10.16=81.28 = 0.8%

OpenStudy (scott208):

@mathmale

OpenStudy (mathmale):

You were to find 125% of 8128. As before, I suggest you convert "125%" into the mixed fraction 1.25 and then multiply 8128 by 1.25. Your "100%=10160" is, unfortunately, incorrect, both in quantity and in terms of math. The following would be correct; 100% of 8128 is simply 8128, obtained by multiplying 8128 by 1. 25% of 8128 is 2032, obtained by multiplying 8128 by 0.25. So 125% of 8128 is obtained by adding these two results together, or by multiplying 8128 by 1.25. That comes out to 10160. Now, find 0.8% of 10160.

OpenStudy (scott208):

i just did